A professor wants to randomly select 4 students to go to the board. She decides to randomly select the fifth

student who enters the classroom and every sixth student after that. Determine the students who will be going to the board.

Write down the student numbers.

1.4.RA-3 0 of 1 Point EQuestion Help A professor wants to randomly select 4 students to go to the board. She decides to randomly select the fifth student who enters the classroom and every sixth student after that. Determine the students who will be going to the board. Write down the student numbers. The numbers of the students that will be going to the board are (Use a comma to separate answers as needed.) Enter your answer in the answer box and then click Check Answer. All parts showing Clear All Check Answer
Answer #1

5th student is selected

5+6 = 11th student is selected

11+6 = 17th student is selected

17+6 = 23rd student is selected

The number of students that will be going to the board are 5,11,17,23.